Grade 12th passAlgebra

if a and b are the roots of ax^2+2bx+c=0,and y,s are the roots of Ax^2+Bx+c=0 such that a,b,y,s are in ap then b^2-ac/B^2-AC is

Sushilkumar
9 Years agoGrade 12th pass

1 Answer

Varun
9 Years ago
I think that there is a mistake in question it should be b^2-ac/B^2-4AC. If question is like that then solution is a+b=-2b/a ,ab=c/a,y+s =-B/A,YS= C/A , as they are in ap a-b =y-s use this you get Answer as a^2/4A^2